# Admissibility of Stein Shrinkage for Batch Normalization in the Presence of Adversarial Attacks

### v1

@@ 6 Acknowledgments
This research was supported in part by NIH National Institute of Neurological Disorders (NINDS) and National Institute for Aging (NIA) under Grant RO1NS121099. (Corresponding author: Baba C. Vemuri.)
@@ Appendix A Technical Appendices and Supplementary Material
In this Appendix, we present the detailed proofs of the two key theorems 3.1 and 3.3 and additional experiments showing Stein’s correction applied only to the mean and not the variance. These experimental results are depicted here in this Appendix as additional rows added in to the three tables appearing in the main body of the paper. Also included are further experimental details.
Theorem 3.1 Proof:.
We denote MLE estimator as $\hat{\theta}^{\mathrm{MLE}}=z=\left(z_{1},\ldots z_{p}\right).$
Then James - Stein Estimator is $\hat{\theta}^{\mathrm{JS}}=z-\frac{(p-2)z}{z^{\top}z}=z\left(1-\frac{p-2}{z^{\top}z}\right)=\hat{\theta}^{\mathrm{MLE}}\left(1-\frac{p-2}{\hat{\theta}^{\mathrm{MLE}\top}\hat{\theta}^{\mathrm{MLE}}}\right).$
The risk of James - Stein Estimator equals to
$R(\hat{\theta}^{\mathrm{JS}},\theta)$ $=\mathbb{E}\left(\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)^{\top}\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)\right)=$
@@
$=R(\hat{\theta}^{\mathrm{MLE}})-2(p-2)\mathbb{E}\left(\frac{z^{\top}(z-\theta)}{z^{\top}z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{z^{\top}z}\right)$
@@
$-2(p-2)\mathbb{E}\left(\frac{z^{\top}(z-\theta)}{z^{\top}z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{z^{\top}z}\right)=$
@@
We want to show that $\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)<2$
Let us denote $\tilde{x}=2z^{\top}\theta+p-2\quad$ and $\tilde{y}=\frac{1}{z^{\top}z}.$
By applying Cauchy–Schwarz inequality we get,
$\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)=\mathbb{E}(\tilde{x}\cdot\tilde{y})\leq|\mathbb{E}(\tilde{x}\cdot\tilde{y})|\leq\sqrt{\mathbb{E}\left(\tilde{x}^{2}\right)}\sqrt{\mathbb{E}\left(\tilde{y}^{2}\right)}.$
Let us calculate each term separately:
$\mathbb{E}\left(\tilde{x}^{2}\right)=\mathbb{E}\left(\left(2z^{\top}\theta+p-2\right)^{2}\right)=\mathbb{E}\left(\left(2z^{\top}\theta\right)^{2}+4\left(z^{\top}\theta\right)(p-2)+(p-2)^{2}\right)=$
$=\mathbb{E}\left(4\left(x^{\top}\theta+y^{\top}\theta\right)^{2}+4\left(x^{\top}\theta+y^{\top}\theta\right)(p-2)\right)+(p-2)^{2}=$
$=4\mathbb{E}\left((x^{\top}\theta)^{2}\right)+8\mathbb{E}\left((x^{\top}\theta)(y^{\top}\theta)\right)+4\mathbb{E}\left((y^{\top}\theta)^{2}\right)+4(p-2)\mathbb{E}\left(x^{\top}\theta\right)+(p-2)^{2}=$
$=4Var\left(x^{\top}\theta\right)+4\left(\mathbb{E}\left(x^{\top}\theta\right)\right)^{2}+4Var\left(y^{\top}\theta\right)+4\left(\mathbb{E}\left(y^{\top}\theta\right)\right)^{2}+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}=$
$=4||\theta||_{2}^{2}+4(||\theta||_{2}^{2})^{2}+4Var\left(y^{\top}\theta\right)+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}\leq$
$4||\theta||_{2}^{2}+4(||\theta||_{2}^{2})^{2}+8\epsilon^{2}||\theta||_{2}^{2}+4(p-2)||\theta||_{2}^{2}+(p-2)^{2}=4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}.$
So, $\sqrt{\mathbb{E}\left(\tilde{x}^{2}\right)}\leq\sqrt{4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}}.$
Note that,
$\mathbb{E}\left(z^{\top}z\right)=\mathbb{E}\left(\sum_{i}z_{i}^{2}\right)=\sum_{i}\mathbb{E}\left(z_{i}^{2}\right)=\sum_{i}Var(z_{i})+\sum_{i}\left(\mathbb{E}(z_{i})\right)^{2}\geq p+||\theta||_{2}^{2}.$
Using Jensen inequality, firstly for $f(x)=\frac{1}{x}$ and then for $g(x)=x^{2}$ we get
$\sqrt{\mathbb{E}\left(\tilde{y}^{2}\right)}=\sqrt{\mathbb{E}\left(\frac{1}{(z^{\top}z)^{2}}\right)}\leq\sqrt{\frac{1}{\mathbb{E}\left((z^{\top}z)^{2}\right)}}\leq\frac{1}{\mathbb{E}\left(z^{\top}z\right)}\leq\frac{1}{p+||\theta||_{2}^{2}}.$
So, overall we get
$\mathbb{E}\left(\frac{2z^{\top}\theta+p-2}{z^{\top}z}\right)\leq\frac{\sqrt{4||\theta||_{2}^{2}\left(||\theta||_{2}^{2}+2\epsilon^{2}+p-1\right)+(p-2)^{2}}}{p+||\theta||_{2}^{2}}\leq 2.$
The final inequality can be established by direct computation. ∎
Theorem 3.3 Proof:.
@@
[…]
@@
    $02\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>$ 0 (7)
@@
    $02\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<$ 0 (8)
Let us derive some facts that will be helpful in the proof.
Firstly, we aim to show that for each $i$, the following inequality holds:
@@
Reminder. A random variable $Z$ is sub-exponential if:
$\mathbb{E}[e^{\lambda Z}]\leq\exp\left(\frac{\lambda^{2}\nu^{2}}{2}\right),\quad|\lambda|<\frac{1}{b}.$
Bernstein Inequality for Sub-Exponential Random Variables. Let $X_{1},\dots,X_{n}$ be independent, centered sub-exponential variables with parameters $(\nu,b)$. Then for any $t>0$:
@@
where $\delta:=\exp(-nC)\to 0$ as $n\to\infty$. Thus, the desired non-negativity condition holds with high probability in the large sample regime. We will use this fact in the proof of case 2. Now we are ready to consider each case separately.
Case 1. We examine the first term in more detail:
@@
Since $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$, where the parameters $\alpha$ and $\beta$ are known, we can apply Stein’s lemma for the Gamma distribution:
Lemma A.1 (Stein’s Lemma for the Gamma Distribution).
Then the following identity holds:
@@
We apply Stein’s identity to the function $h(x)=V$, where
$V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+w_{i}\right)\right]^{1/p},\quad X_{i}=\hat{\sigma}_{x,i}^{2}.$
Using the Gamma Stein identity:
we differentiate $h(X)$ with respect to $\hat{\sigma}_{x,k}^{2}$ as follows:
@@
Substituting into Stein’s identity for $X_{i}=\hat{\sigma}_{x,i}^{2}$, we get:
In particular, we obtain the inequality:
$\mathbb{E}[V]\leq\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]\quad\Rightarrow\quad-\mathbb{E}[V]\geq-\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right].$
Using this bound equation (9) becomes:
@@
$=2\sum_{i=1}^{p}\left[\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}V\right]\left(\frac{1}{\alpha+1}-\frac{1}{\alpha}\right)+\frac{1}{\alpha+1}\mathbb{E}\left[\hat{\sigma}_{y,i}^{2}V\right]+\frac{1}{\alpha+1}\mathbb{E}\left[W_{i}V\right]\right]\geq$
$2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}\right)V\right]\cdot\left(-\frac{1}{\alpha(\alpha+1)}\right)\right)$
Let’s denote
$\overline{\hat{\sigma}_{z}^{2}}=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2}.$
Overall, we get lower bound for :
@@
Last inequality appears due to assumption. So we have:
$c<0\ \text{and}\ \frac{2}{\alpha(\alpha+1)}+c<0\Rightarrow c\leq-\frac{2}{\alpha(\alpha+1)}.$
@@
$02\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)+pc\,\mathbb{E}(V^{2})<$ 0 .
Recall that using Bernstein’s inequality we get with high probability, the term $\hat{\sigma}_{y,i}^{2}+W_{i}$ is non-negative. So, combining this with Stein’s lemma for $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$ we get:
@@
$\geq\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)-\frac{1}{\alpha p}\mathbb{E}(V)$
$\mathbb{E}(V)\left(1+\frac{1}{\alpha p}\right)$ $\geq\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)$
@@
Plug in this bound and get inequality for  (9):
@@
$\leq 2\sum_{i=1}^{p}\left(\mathbb{E}\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}\cdot V\right)-\frac{\beta_{i}}{\alpha\beta_{i}}\cdot\frac{1}{1+\frac{1}{\alpha p}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)+\mathbb{E}\left(\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}\cdot V\right)\right)$
Using our previous notation $\overline{\hat{\sigma}_{z}^{2}}=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2},$
$=\frac{2p}{\alpha+1}\mathbb{E}\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)-\frac{2}{\alpha(1+\frac{1}{\alpha p})}\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right).$
Overall for (8) we get:
@@
The last inequality we want to establish by finding the appropriate $c$. So, we will rewrite it in a more suitable way:
@@
Let’s bound each term independently. Note that $\overline{\hat{\sigma}_{z}^{2}}\geq V$, then we bound second term
@@
$\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)=\left|\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)\right|\leq\sqrt{\mathbb{E}((\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2})^{2})}\cdot\sqrt{\mathbb{E}(V^{2})}.$
For the first term we get:
@@
By applying Jensen’s inequality for $f(x)=\sqrt{x}$ – concave, we have:

### v2

@@ 5 Discussion and Conclusion
This research was supported in part by NIH National Institute of Neurological Disorders (NINDS) and National Institute for Aging (NIA) under Grant RO1NS121099 to Vemuri. It was also partially supported by the NSF Smart and Connected Health grant 2205417 to Fletcher.
PPMI Data Availability Statement: Data used in the preparation of this article were obtained from the Parkinson’s Progression Markers Initiative (PPMI) database [(https://www.ppmi-info.org/access-dataspecimens/download-data, accessed on 6 December 2023, RRID:SCR 006431]. For up-to-date information on the study, visit www.ppmi-info.org. This analysis used data openly available from PPMI, obtained from PPMI upon request.
PPMI Data Acknowledgments: We extend our gratitude to Lesa Melnyczuk and Iryna Milne for their valuable assistance. Data used in the preparation of this article were obtained on 6 December, 2023 from the Parkinson’s Progression Markers Initiative (PPMI) database (https://www.ppmi-info.org/access-data-specimens/download-data, accessed on 6 Decemeber 2023), RRID:SCR-006431. For up-to-date information on the study, visit http://www.ppmi-info.org, accessed on 6 Dec. 2023. PPMI—a public–private partnership—is funded by the Michael J. Fox Foundation for Parkinson’s Research and funding partners, including 4D Pharma, Abbvie, AcureX, Allergan, Amathus Therapeutics, Aligning Science Across Parkinson’s, AskBio, Avid Radiopharmaceuticals, BIAL, BioArctic, Biogen, Biohaven, BioLegend, BlueRock Therapeutics, Bristol-Myers Squibb, Calico Labs, Capsida Biotherapeutics, Celgene, Cerevel Therapeutics, Coave Therapeutics, DaCapo Brainscience, Denali, Edmond J. Safra Foundation, Eli Lilly, Gain Therapeutics, GE HealthCare, Genentech, GSK, Golub Capital, Handl Therapeutics, Insitro, Jazz Pharmaceuticals, Johnson & Johnson Innovative Medicine, Lundbeck, Merck, Meso Scale Discovery, Mission Therapeutics, Neurocrine Biosciences, Neuron23, Neuropore, Pfizer, Piramal, Prevail Therapeutics, Roche, Sanofi, Servier, Sun Pharma Advanced Research Company, Takeda, Teva, UCB, Vanqua Bio, Verily, Voyager Therapeutics, the Weston Family Foundation, and Yumanity Therapeutics.
@@ Appendix A Technical Appendices and Supplementary Material
In this Appendix, we present the detailed proofs of the two key theorems 1 and 3, and provide a detailed discussion of the experimental setups.
Proof [Theorem 1 Proof:]
We denote the observation-based estimator by
$\hat{\theta}^{\mathrm{0}}=Z=(Z_{1},\ldots,Z_{p}).$
Then the James–Stein estimator is
$\hat{\theta}^{\mathrm{JS}}=Z-\frac{(p-2)Z}{Z^{\top}Z}=Z\left(1-\frac{p-2}{Z^{\top}Z}\right)=\hat{\theta}^{\mathrm{0}}\left(1-\frac{p-2}{\hat{\theta}^{\mathrm{0}\top}\hat{\theta}^{\mathrm{0}}}\right).$
The risk of James - Stein Estimator is given by,
$R(\hat{\theta}^{\mathrm{JS}},\theta)$ $=\mathbb{E}\left(\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)^{\top}\left(\hat{\theta}^{\mathrm{JS}}-\theta\right)\right)=$ (7)
@@
$=R(\hat{\theta}^{\mathrm{0}})-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right)$
@@
$-2(p-2)\mathbb{E}\left(\frac{Z^{\top}(Z-\theta)}{Z^{\top}Z}\right)+(p-2)^{2}\mathbb{E}\left(\frac{1}{Z^{\top}Z}\right)=$ (8)
@@
We want to show that
$\mathbb{E}\!\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)<2.$
For convenience, set
$\tilde{x}=2Z^{\top}\theta+p-2,\quad\tilde{y}=\frac{1}{Z^{\top}Z}.$
Then,
[…]
By the Cauchy–Schwarz inequality,
$\mathbb{E}(\tilde{x}\,\tilde{y})\leq\sqrt{\mathbb{E}(\tilde{x}^{2})}\,\sqrt{\mathbb{E}(\tilde{y}^{2})}.$
Let us calculate each term separately. First,
$\mathbb{E}\!\left(\tilde{x}^{2}\right)$ $=\mathbb{E}\!\left(\big(2Z^{\top}\theta+p-2\big)^{2}\right)$ (9)
$=\mathbb{E}\!\left((2Z^{\top}\theta)^{2}+4(p-2)(Z^{\top}\theta)+(p-2)^{2}\right)$
$=4\,\mathbb{E}\!\left((Z^{\top}\theta)^{2}\right)+4(p-2)\,\mathbb{E}(Z^{\top}\theta)+(p-2)^{2}.$
Recall that $Z=X+Y$, where $X\sim\mathcal{N}(\theta,I_{p})$ with mean $\theta$ and $Y$ is a mean-zero sub-Gaussian noise. Hence
$\mathbb{E}\!\left((Z^{\top}\theta)^{2}\right)$ $=\mathbb{E}\!\left((X^{\top}\theta+Y^{\top}\theta)^{2}\right)$ (10)
$=\mathbb{E}\!\left((X^{\top}\theta)^{2}\right)+2\,\mathbb{E}\!\left((X^{\top}\theta)(Y^{\top}\theta)\right)+\mathbb{E}\!\left((Y^{\top}\theta)^{2}\right).$
Now
$\mathbb{E}\!\left((X^{\top}\theta)^{2}\right)=\operatorname{Var}(X^{\top}\theta)+\big(\mathbb{E}(X^{\top}\theta)\big)^{2}=\|\theta\|_{2}^{2}+\|\theta\|_{2}^{4},$ (11)
and
$\mathbb{E}(X^{\top}\theta)=\|\theta\|_{2}^{2}.$ (12)
For the noise term, by sub-Gaussianity we get,
$\mathbb{E}\!\left((Y^{\top}\theta)^{2}\right)\leq 2\varepsilon^{2}\,\|\theta\|_{2}^{2}.$ (13)
Substituting (10)–(13) into (9) we obtain,
$\mathbb{E}\!\left(\tilde{x}^{2}\right)\leq 4\|\theta\|_{2}^{2}\left(\|\theta\|_{2}^{2}+2\varepsilon^{2}+p-1\right)+(p-2)^{2}.$
Therefore,
$\sqrt{\mathbb{E}\!\left(\tilde{x}^{2}\right)}\;\leq\;\sqrt{\,4\|\theta\|_{2}^{2}\big(\|\theta\|_{2}^{2}+2\varepsilon^{2}+p-1\big)+(p-2)^{2}}.$ (14)
First, observe that
$\mathbb{E}\!\left(Z^{\top}Z\right)=\sum_{i=1}^{p}\mathbb{E}(Z_{i}^{2})=\sum_{i=1}^{p}\operatorname{Var}(Z_{i})+\sum_{i=1}^{p}\big(\mathbb{E}(Z_{i})\big)^{2}\;\;\geq\;p+\|\theta\|_{2}^{2}.$
Next, by Jensen’s inequality applied to the convex functions $f(x)=1/x$ and $g(x)=x^{2}$, we have
$\sqrt{\mathbb{E}\!\left(\tilde{y}^{2}\right)}$ $=\sqrt{\mathbb{E}\!\left(\frac{1}{(Z^{\top}Z)^{2}}\right)}$ (15)
$\leq\sqrt{\frac{1}{\mathbb{E}\!\left((Z^{\top}Z)^{2}\right)}}\;\leq\;\frac{1}{\mathbb{E}(Z^{\top}Z)}$
$\leq\frac{1}{p+\|\theta\|_{2}^{2}}.$
Combining (14) and (15), we obtain
$\mathbb{E}\!\left(\frac{2Z^{\top}\theta+p-2}{Z^{\top}Z}\right)\;\leq\;\frac{\sqrt{\,4\|\theta\|_{2}^{2}\big(\|\theta\|_{2}^{2}+2\varepsilon^{2}+p-1\big)+(p-2)^{2}}}{p+\|\theta\|_{2}^{2}}\;\leq\;2.$
The last inequality can be verified by direct computation.
Proof [Theorem 3 Proof:]
@@
$R(\hat{\beta}^{\,\mathrm{JS}},\beta)=\mathbb{E}\left(\left\|\hat{\beta}^{\,\mathrm{JS}}-\beta\right\|_{2}^{2}\right)=\mathbb{E}\left(\left\|\hat{\beta}^{\,0}+cV-\beta\right\|_{2}^{2}\right)=$
$=\mathbb{E}\left(\left\|\hat{\beta}^{\,0}-\beta\right\|_{2}^{2}\right)+2\sum_{i=1}^{p}\mathbb{E}\left(\left(\hat{\beta}_{0,i}-\beta_{i}\right)\cdot cV\right)+p\mathbb{E}\left(c^{2}V^{2}\right)$
@@
    $2\sum_{i=1}^{D}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]>0$ (16)
@@
    $2\sum_{i=1}^{p}\mathbb{E}\left[\left(\hat{\beta}_{i}^{0}-\beta_{i}\right)V\right]+pc\,\mathbb{E}\left[V^{2}\right]<0$ (17)
We now establish an intermediate result in the form of a lemma that will be useful in the proof of Theorem 3.
Lemma 6 (Probability bound for $\hat{\sigma}^{2}_{y,i}+W_{i}$)
Let $X_{ij}\sim\mathcal{N}(\mu,\sigma_{x,i}^{2})$ and $Y_{ij}\sim\mathrm{SG}(2\varepsilon^{2})$ be independent for $j=1,\dots,n$. Define
$\hat{\sigma}_{y,i}^{2}=\frac{1}{n}\sum_{j=1}^{n}(Y_{ij}-\bar{Y}_{i})^{2},\quad W_{i}=\frac{2}{n}\sum_{j=1}^{n}(X_{ij}-\bar{X}_{i})(Y_{ij}-\bar{Y}_{i}),$
where $\bar{X}_{i}=\frac{1}{n}\sum_{j}X_{ij}$, $\bar{Y}_{i}=\frac{1}{n}\sum_{j}Y_{ij}$. Then, for any $\delta\in(0,1)$,
@@
We now recall the Bernstein inequality that will be used to prove the desired bound. Bernstein Inequality for Sub-Exponential Random Variables: Let $X_{1},\dots,X_{n}$ be independent, centered sub-exponential [^1] A random variable $Z$ is sub-exponential with parameters $(\nu,b)$ if $\mathbb{E}[e^{\lambda Z}]\leq\exp(\lambda^{2}\nu^{2}/2)$ for all $|\lambda|<1/b$. variables with parameters $(\nu,b)$. Then for any $t>0$:
@@
where $\delta:=\exp(-nC)\to 0$ as $n\to\infty$. Thus, the desired non-negativity condition holds with high probability in the large sample regime.
We will use this lemma in the proof of case 2. Now we are ready to consider each case separately.
Case 1. We examine the first term of (16) in more detail:
@@
Since $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$, where the parameters $\alpha$ and $\beta$ are known, we can now apply Stein’s Lemma for the Gamma Distribution, which states: Let $X\sim\mathrm{Gamma}(\alpha,\beta)$, and let $h(x)$ be a differentiable function such that $\mathbb{E}[Xh(X)]<\infty$ and $\mathbb{E}[Xh^{\prime}(X)]<\infty.$ Then the following identity holds:
@@
This can be rewritten as:
$\mathbb{E}[h(X)]=\frac{1}{\alpha\beta}\mathbb{E}[Xh(X)]-\frac{1}{\alpha}\mathbb{E}[Xh^{\prime}(X)].$ (19)
We now apply Stein’s identity in (19) to the function $h(X)=V$, where
$V=\left[\prod_{i=1}^{p}\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}\right)\right]^{1/p}.$
The derivative of $V$ with respect to $\hat{\sigma}_{x,k}^{2}$ is:
@@
Substituting Stein’s identity (19) for $\hat{\sigma}_{x,i}^{2}$, we get the following.
$\mathbb{E}[V]=\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]-\frac{1}{\alpha p}\mathbb{E}\left[V\cdot\frac{\hat{\sigma}_{x,i}^{2}}{\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}}\right].$ (20)
In particular, we can obtain the following inequality:
$\mathbb{E}[V]\leq\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right]\quad\Rightarrow\quad-\mathbb{E}[V]\geq-\frac{1}{\alpha\beta_{i}}\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}\cdot V\right].$ (21)
Using the above bound in (21), equation (18) becomes:
@@
$2\sum_{i=1}^{p}\left[\mathbb{E}\left[\hat{\sigma}_{x,i}^{2}V\right]\left(\frac{1}{\alpha+1}-\frac{1}{\alpha}\right)+\frac{1}{\alpha+1}\mathbb{E}\left[\hat{\sigma}_{y,i}^{2}V\right]+\frac{1}{\alpha+1}\mathbb{E}\left[W_{i}V\right]\right]\geq$
$2\sum_{i=1}^{p}\left(\mathbb{E}\left[\left(\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}\right)V\right]\cdot\left(-\frac{1}{\alpha(\alpha+1)}\right)\right)$ (22)
Recall that in the statement of the theorem we introduced $\hat{\sigma}_{z,i}^{2}=\hat{\sigma}_{x,i}^{2}+\hat{\sigma}_{y,i}^{2}+W_{i}$. For convenience, in the proof we will work with the averaged version
$\overline{\hat{\sigma}_{z}^{2}}=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2}.$ (23)
From (22) the lower bound for (16) is given by:
@@
The last inequality follows from our case-1 hypothesis. Therefore,
[…]
@@
$2\sum_{i=1}^{p}\mathbb{E}\left((\hat{\beta}_{0,i}-\beta_{i})V\right)+pc\,\mathbb{E}(V^{2})<0.$
Recall that using the lemma 6, with high probability, we find that the term $\hat{\sigma}_{y,i}^{2}+W_{i}$ is non-negative. So, combining this with Stein’s lemma (19) applied to $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta)$ and $V$ we get:
@@
$\geq\frac{1}{\alpha\beta_{i}}\,\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)-\frac{1}{\alpha p}\,\mathbb{E}(V).$
$\mathbb{E}(V)\left(1+\frac{1}{\alpha p}\right)$ $\geq\frac{1}{\alpha\beta_{i}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)$ (24)
@@
Substituting the bound (24) into (18), we get the following inequality:
@@
$\leq 2\sum_{i=1}^{p}\left(\mathbb{E}\left(\frac{\hat{\sigma}_{x,i}^{2}}{\alpha+1}\cdot V\right)-\frac{\beta_{i}}{\alpha\beta_{i}}\cdot\frac{1}{1+\frac{1}{\alpha p}}\mathbb{E}(\hat{\sigma}_{x,i}^{2}\cdot V)+\mathbb{E}\left(\frac{\hat{\sigma}_{y,i}^{2}+W_{i}}{\alpha+1}\cdot V\right)\right)$ (25)
As defined in (23) $\overline{\hat{\sigma}_{z}^{2}}:=\frac{1}{p}\sum_{i=1}^{p}\hat{\sigma}_{z,i}^{2}$ , we can rewrite the expression (25) as
$\frac{2p}{\alpha+1}\,\mathbb{E}\!\left(\overline{\hat{\sigma}_{z}^{2}}\cdot V\right)-\frac{2}{\alpha\left(1+\frac{1}{\alpha p}\right)}\,\mathbb{E}\!\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right).$ (26)
So, for (17) we get the following bound by applying (26):
@@
By finding the appropriate $c$ we can obtain the last inequality.
Rewriting it in a more cleaner way, we have:
@@
Let us now bound each term (on the RHS of the above inequality) independently. Note that $\overline{\hat{\sigma}_{z}^{2}}\geq V$, so we can write a bound for second term in (27)
@@
$\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)=\left|\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)\right|\leq\sqrt{\mathbb{E}((\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2})^{2})}\cdot\sqrt{\mathbb{E}(V^{2})}.$ (29)
Using (29) we can bound the first term in (27):
@@
By applying Jensen’s inequality to the function of the form: $f(x)=\sqrt{x}$, we have: